# A Digital Twin-Driven Recommendation System for Adaptive Campus Course Timetabling

### v1

@@ Abstract
Efficient and adaptive course timetabling for large, dynamic university campuses remains a significant challenge due to the complex interplay of hard and soft constraints. Traditional static optimization methods often fail to accommodate real-time disruptions, evolving user preferences, and the nuanced spatial–temporal relationships inherent in campus environments. This paper reconceptualizes the timetabling problem as a recommendation-based task and leverages the Texas A&M Campus Digital Twin as a dynamic data platform. Our proposed framework integrates collaborative and content-based filtering techniques with iterative feedback mechanisms, thereby generating a ranked set of adaptive timetable recommendations. A composite scoring function, incorporating metrics for classroom occupancy, travel distance, travel time, and vertical transitions, enables the framework to systematically balance resource utilization with user-centric factors. Extensive experiments using real-world data from Texas A&M University demonstrate that our approach effectively reduces travel inefficiencies, optimizes classroom utilization, and enhances overall user satisfaction. By coupling a recommendation-oriented paradigm with a digital twin environment, this study offers a robust and scalable blueprint for intelligent campus planning and resource allocation, with potential applications in broader urban contexts.
@@
Course Timetabling , Recommendation Systems , Digital Twin , Spatial-Temporal Analysis , Iterative Feedback , Campus Planning
@@
[label1]organization=Department of Landscape Architecture & Urban Planning and The Center for Geospatial Sciences, Applications, & Technology, Texas A&M University,city=College Station, state=TX, country=USA
@@
[label2]organization=Department of Geography & Environmental Sustainability, The University of Oklahoma,city=Norman, state=OK, country=USA
@@ 1 Introduction
Timetabling in large academic institutions represents a multifaceted challenge that requires the precise allocation of courses, classrooms, and time slots while meeting rigorous institutional policies and addressing the diverse needs of students and faculty [1, 2, 3]. The problem is compounded by the necessity to balance hard constraints—such as classroom capacities, scheduling conflicts, and feasible travel times—with soft constraints including faculty preferences, student satisfaction, and inter-departmental coordination [4, 5]. Furthermore, spatial–temporal factors such as building locations, travel routes, and transition times add a critical dimension to the problem. These elements, along with dynamic influences like fluctuating enrollment numbers and changing infrastructure conditions, render traditional static optimization approaches insufficient [6, 7, 8, 9]. Consequently, there is an urgent need for innovative, adaptive methodologies that comprehensively integrate spatial, temporal, and operational considerations.
@@
Extensive research has addressed the complexities of course timetabling using a variety of optimization techniques. Foundational approaches—such as constraint satisfaction models [10] and combinatorial optimization frameworks [11]—have laid the groundwork for automated timetabling. In addition, evolutionary algorithms, including genetic algorithms [12], harmony search [13], and particle swarm optimization [14], have demonstrated the ability to navigate large solution spaces and balance competing objectives effectively. Mixed-integer programming (MIP) techniques provide exact solutions for smaller problems [15, 16, 17], while multi-objective optimization methods adeptly manage trade-offs between conflicting goals, such as minimizing travel distances while maximizing classroom utilization [18]. Moreover, hyper-heuristics have emerged as a flexible means to address diverse scheduling scenarios by balancing computational efficiency and solution quality [11].
@@
Recent advancements have further enhanced traditional methods by integrating machine learning and hybridization strategies. For instance, fuzzy logic has enabled the dynamic incorporation of user preferences into the scheduling process [19], and machine learning techniques have improved algorithm selection and design [20, 21]. Clustering algorithms have been employed to simplify complex schedules [22], and reinforcement learning has demonstrated promise in refining schedules through iterative interactions with dynamic environments [23, 24]. Despite these advancements, significant gaps remain. Many existing approaches do not fully incorporate spatial–temporal correlations or multi-modal transportation options—factors that are critical for optimizing travel efficiency and resource allocation in large academic campuses. Additionally, most models lack the capability to generate adaptive schedules that respond in real time to disruptions and evolving requirements.
@@
Digital twin technology has recently emerged as a transformative tool for addressing these challenges [25, 26]. A digital twin is a dynamic, real-time virtual representation of a physical system that seamlessly integrates spatial–temporal data with operational constraints, offering a powerful platform for simulation, monitoring, and optimization. While digital twins have been widely adopted in manufacturing to enhance production scheduling and operational efficiency [27, 28, 29], their application in urban systems and academic environments is still emerging. Recent work in academic scheduling demonstrates the potential of digital twins to adapt dynamically to changes such as fluctuating enrollments, evolving classroom availability, and infrastructure disruptions [25]. By merging diverse datasets—including classroom attributes, transportation networks, and user feedback—digital twin technology enables a holistic approach to optimizing campus operations.
@@
In this study, we leverage the Texas A&M Campus Digital Twin to bridge the identified research gaps. We propose an iterative optimization framework for course timetabling that integrates spatial–temporal correlations, multi-modal transportation considerations, and user-centric metrics. A composite scoring function is used to evaluate course assignments based on normalized measures of classroom occupancy, travel distance, travel time, and vertical transitions between floors. The iterative optimization process generates multiple feasible alternatives for bottleneck courses, ensuring targeted adjustments and dynamic responsiveness. By incorporating real-time feedback and leveraging digital twin capabilities, the proposed framework offers a robust and scalable solution for the complex, dynamic scheduling requirements of large academic environments.
@@
This research makes several key contributions. First, it integrates spatial–temporal factors and multi-modal transportation options into the course scheduling process, addressing a critical gap in existing methodologies. Second, it demonstrates the value of an iterative, feedback-driven optimization process for dynamically adapting to evolving constraints and user needs. Third, it highlights the broader applicability of this approach beyond academic scheduling, extending to urban resource management and public facility planning. By emphasizing adaptability, transparency, and user-centricity, our work provides a comprehensive framework for addressing the complexities of modern timetabling.
@@ 2.1 Problem Settings
Large-scale course timetabling at universities such as Texas A&M University involves the systematic assignment of a set of courses $C=\{c_{i}\mid i=1,\ldots,N\}$ to a network of classrooms $R=\{r_{j}\mid j=1,\ldots,M\}$ across a series of time slots $T=\{t_{k}\mid k=1,\ldots,K\}$ for a student population $V=\{v_{l}\mid l=1,\ldots,L\}$. The objective is to construct a timetable that not only satisfies stringent institutional policies but also aligns with user preferences and optimizes resource utilization. The complexity arises from the need to meet both hard constraints (e.g., classroom capacity, conflict avoidance, and travel feasibility) and soft constraints (e.g., faculty preferences, degree requirements, and inter-departmental coordination). Moreover, spatial–temporal factors—such as building locations, travel routes, and transition times—introduce additional challenges. This multifaceted problem thus requires a comprehensive framework that assigns courses while simultaneously optimizing the ease of student movement across campus.
@@ 2.2 Variables and Data Representation
    $r^{\text{prev}}(v_{l},t_{k})=\begin{cases}r_{j^{\prime}}&\text{if }\exists\,c_{i^{\prime}}\text{ and }t_{k-1}\text{ such that }x_{i^{\prime}j^{\prime}(k-1)}=1\text{ and }v_{l}\in\phi(c_{i^{\prime}}),\\ \varnothing&\text{otherwise,}\end{cases}$ (1)
@@
Spatial–Temporal Data
@@
$x_{ijk}=\begin{cases}1,&\text{if course }c_{i}\text{ is assigned to classroom }r_{j}\text{ at time }t_{k},\\ 0,&\text{otherwise.}\end{cases}$
@@ 2.3 Objective Function
Maximize $\text{Maximize }Z=\alpha_{1}\bar{S}_{O}+\alpha_{2}\bar{S}_{D}+\alpha_{3}\bar{S}_{T}+\alpha_{4}\bar{S}_{F},$ (10)
@@ 2.6 Complexity and Proposed Solution
The integration of numerous hard and soft constraints renders the course timetabling problem a high-dimensional, combinatorial challenge. Traditional static optimization methods often struggle to capture the dynamic, user-centric aspects inherent in such problems. To address these challenges, we propose an iterative recommendation framework that incorporates real-time feedback, detailed spatial–temporal analysis, and dynamic scoring adjustments. This adaptive approach is designed to converge to a high-quality timetable that effectively harmonizes institutional objectives with user satisfaction, thus achieving enhanced adaptability, robustness, and operational efficiency.
@@ 3.3 System Highlights
The proposed recommendation system leverages both spatial–temporal factors and multi-modal transportation data to produce high-quality timetabling solutions. Two key aspects of the system are highlighted below, each incorporating quantitative measures to enhance decision-making.
@@ 4.3 Metric-Based Evaluation
Overall, these metric-based assessments demonstrate the system’s capacity to generate practical, data-driven recommendations that support the twin goals of flexibility and user-centered optimization. By quantifying changes in travel distance, travel time, occupancy, and floor transitions, the framework ensures that proposed reassignments systematically address previously identified spatial–temporal bottlenecks, thereby improving the robustness and efficiency of the course timetable.
@@ 4.4 Comparison of Original and Optimized Schedules
In the histograms on the left, the optimized schedule shows a pronounced shift toward lower travel distances and shorter travel times. The reduction in outliers—students or faculty facing especially long commutes—highlights the system’s effectiveness in targeting spatial–temporal bottlenecks identified during earlier analyses. Additionally, the histogram of floors traveled reveals fewer high-floor transitions in the optimized schedule, suggesting that reassignments have reduced vertical movements between classes. Meanwhile, the occupancy rate distribution indicates a slightly more balanced utilization of classrooms.
@@
Collectively, these results demonstrate how iterative feedback loops—guided by the composite scoring function—can systematically refine scheduling decisions. By quantifying improvements in travel distance, travel time, floors traveled, and occupancy, the framework ensures that adjustments made in each iteration yield tangible benefits. The final outcome is a more user-centric timetable that balances resource utilization with spatial–temporal feasibility, reflecting the adaptive nature of the proposed methodology.
@@ 5 Discussion
A key insight from these experiments was the reduction in travel-related inefficiencies for students and faculty. By incorporating spatial–temporal data, the framework pinpointed and remedied courses that contributed disproportionately to lengthy transitions between classes. In practical terms, this meant fewer instances of extended walks or rushed travel times, which in turn fostered an improved sense of schedule convenience among users. Furthermore, the system’s iterative approach enabled it to react quickly to unexpected events by refining course assignments incrementally, rather than recalculating the entire timetable. Another notable outcome was the framework’s capacity to optimize classroom utilization while respecting critical constraints such as room capacity and scheduling conflicts. By reallocating courses to underused spaces during busy periods, the system minimized congestion without requiring a full overhaul of existing assignments. This flexibility illustrates the framework’s user-centric ethos, wherein data-driven insights guide reassignments that benefit both institutional objectives and individual needs.
@@
Overall, these findings suggest that the proposed recommendation-based framework offers a robust and scalable solution to the complex challenge of course timetabling. By coupling spatial–temporal analytics with iterative refinement and user-centric design, the system not only enhances operational efficiency but also aligns scheduling decisions with the evolving needs of a diverse campus community.
@@ 6 Conclusion
This study introduced an adaptive recommendation-based framework for course timetabling, underpinned by the Texas A&M Campus Digital Twin and designed to accommodate spatial–temporal constraints, evolving user needs, and real-time disruptions. In shifting from a static to a dynamic scheduling paradigm, the framework merges collaborative and content-based filtering techniques, enabling iterative refinement of schedules and ensuring their continued feasibility and efficiency.
The experimental findings confirm the framework’s potential to reduce travel distances and times, enhance classroom occupancy, and improve user satisfaction. By continuously reallocating resources and resolving critical bottlenecks, the system adeptly balances institutional requirements—such as room capacity and scheduling constraints—with student and faculty preferences. Moreover, the framework’s demonstrated scalability suggests that it can be extended to larger and more complex domains, including urban resource management and multimodal transportation networks.
Despite these strengths, the framework relies on high-quality, comprehensive data for optimal performance, highlighting the importance of robust data integration and validation. Incomplete or inaccurate data could undermine its accuracy and scalability. Additionally, while iterative updates are computationally efficient, extreme data volumes or high-frequency changes may necessitate further algorithmic refinements.
Future work could explore the integration of advanced predictive models—potentially incorporating machine learning techniques—to enhance the system’s capacity for forecasting user preferences and preempting dynamic constraints. Expanding the framework to incorporate broader policy objectives, such as equitable resource distribution and sustainability considerations, would also enrich its applicability. By addressing these avenues for improvement, the proposed framework offers a strong foundation for developing scalable, adaptive solutions in both academic timetabling and other large-scale resource management contexts.

### v2

# Human Mobility Reimagined: Digital Twin Intelligence for Adaptive Campus Course Timetabling

@@ Abstract
Daily operations in large campuses depend on how efficiently people move through space and time. In this sense, course timetables are more than administrative schedules: they act as mobility policies that orchestrate thousands of trajectories, shaping travel burden, congestion, accessibility, and the reliability of back-to-back transitions. Designing timetables that are both feasible and mobility-friendly is challenging because hard constraints including capacity, conflicts, feasibility must be satisfied alongside soft constraints including preferences, satisfaction, coordination, all under dynamic conditions such as real-time disruptions and evolving demand. Traditional static optimization methods often struggle to capture these human mobility impacts and to adapt when campus conditions change. This paper reconceptualizes course timetabling as a recommendation-based task and leverages the Texas A&M Campus Digital Twin as a dynamic data platform to evaluate mobility consequences at scale. We propose an iterative framework that integrates collaborative and content-based filtering with feedback-driven refinement to generate ranked sets of adaptive timetable recommendations. A mobility-aware composite scoring function combining classroom occupancy, travel distance, travel time, and vertical transitions systematically balances resource efficiency with human-centered movement costs. Extensive experiments using real-world data from Texas A&M University show that the proposed approach reduces mobility friction and travel inefficiencies, improves classroom utilization, and enhances overall user satisfaction. By coupling recommendation-oriented decision-making with digital twin intelligence, this study provides a robust and scalable blueprint for mobility-centered campus planning and resource allocation, with potential extensions to broader urban systems.
@@
Human Mobility , Course Timetabling , Digital Twin , Spatial-Temporal Analysis , Iterative Feedback , Recommendation Systems
@@
[label1]organization=The Center for Geospatial Sciences, Applications, & Technology, Department of Landscape Architecture and Urban Planning, Texas A&M University,city=College Station, state=TX, country=USA
@@
[label4]organization=Zachry Department of Civil and Environmental Engineering, Texas A&M University,city=College Station, state=TX, country=USA
[…]
[label2]organization=Department of Geography & the Environment, The University of Alabama,city=Tuscaloosa, state=AL, country=USA
[…]
[label5]organization=Texas A&M Institute of Data Science ,city=College Station, state=TX, country=USA
[…]
[label3]organization=Department of Geography & Environmental Sustainability, The University of Oklahoma,city=Norman, state=OK, country=USA
@@ 1 Introduction
Human mobility is a foundational layer of how complex systems function: it determines how people access services, how shared resources are utilized, and how daily activities unfold in space and time [barbosa2018human, gonzalez2008understanding, alessandretti2020scales, huang2025urban]. In increasingly data-rich environments, the central challenge is no longer merely to plan mobility under fixed assumptions, but to adapt mobility-centric decisions as conditions, preferences, and constraints evolve [pappalardo2023future, xugenerative]. Digital twins enable this shift by coupling a physical system with a continuously updated virtual counterpart, making it possible to evaluate decisions not only by feasibility and efficiency, but also by their mobility consequences [aghaabbasisustainability, ye2025toward, lin2024human, wang2022mobility]. This paper adopts that perspective and studies adaptive resource allocation through a human-mobility-first lens, using a university campus as a living microcosm where dense daily movements, tight time budgets, and heterogeneous indoor-outdoor infrastructure make mobility impacts immediately tangible.
@@
Within this setting, a timetable is not simply an administrative artifact; it acts as a mobility policy that orchestrates thousands of day-to-day trajectories. Each assignment of an activity to a time and place implicitly shapes transition pressures, corridor congestion, and the uneven burdens of long walks, delayed arrivals, and vertical movements. Course timetabling in large academic institutions therefore represents a multifaceted challenge that requires the precise allocation of courses, classrooms, and time slots while meeting rigorous institutional policies and addressing the diverse needs of students and faculty [chen2021survey, dinkel1989or, badri1998multi]. The problem is compounded by the necessity to balance hard constraints—such as classroom capacities, scheduling conflicts, and feasible travel times—with soft constraints including faculty preferences, student satisfaction, and inter-departmental coordination [hossain2019optimization, shiau2011hybrid]. Crucially, spatial-temporal factors such as building locations, pedestrian pathways, indoor circulation (e.g., corridors, stairs, elevators), and transition times introduce nontrivial dependencies between consecutive classes. These dependencies vary across individuals, time-of-day conditions, and infrastructure availability, making mobility inherently dynamic. Along with fluctuating enrollment, changing classroom availability, construction, and disruptions, these mobility considerations render traditional static optimization approaches insufficient [burke2002recent, schaerf1999survey, lewis2008survey, hekmati2021course]. Consequently, there is an urgent need for adaptive methodologies that explicitly integrate spatial, temporal, and operational considerations to reduce mobility friction while maintaining institutional feasibility.
@@
Extensive research has addressed the complexities of course timetabling using a variety of optimization techniques. Foundational approaches—such as constraint satisfaction models [mccollum2006perspective] and combinatorial optimization frameworks [qu2009hybridizations]—have laid the groundwork for automated timetabling. In addition, evolutionary algorithms, including genetic algorithms [abdullah2008generating], harmony search [al2012harmony], and particle swarm optimization [tassopoulos2012solving], have demonstrated the ability to navigate large solution spaces and balance competing objectives effectively. Mixed-integer programming (MIP) techniques provide exact solutions for smaller problems [daskalaki2004integer, rappos2022mixed, al2007mixed], while multi-objective optimization methods adeptly manage trade-offs between conflicting goals, such as minimizing travel distances while maximizing classroom utilization [deb2011multi]. Moreover, hyper-heuristics have emerged as a flexible means to address diverse scheduling scenarios by balancing computational efficiency and solution quality [qu2009hybridizations]. Despite these advances, many timetabling formulations still treat mobility as a coarse add-on (e.g., a single distance/time penalty) rather than modeling the full movement experience that includes inter-building flows, indoor-outdoor transitions, vertical movements, and time-varying congestion.
@@
Recent advancements have further enhanced traditional methods by integrating machine learning and hybridization strategies. For instance, fuzzy logic has enabled the dynamic incorporation of user preferences into the scheduling process [chaudhuri2010fuzzy], and machine learning techniques have improved algorithm selection and design [de2022algorithm, soria2014effective]. Clustering algorithms have been employed to simplify complex schedules [kenekayoro2020incorporating], and reinforcement learning has demonstrated promise in refining schedules through iterative interactions with dynamic environments [yue2023reinforcement, obit2011non]. Nevertheless, significant gaps remain from a human mobility perspective. Many existing approaches do not fully incorporate spatial-temporal correlations in real campus networks or account for multi-modal transportation options (walking, bicycling, e-scooters, shuttles) and indoor mobility modes (corridors, stairs, elevators)—factors that directly determine travel feasibility, comfort, and schedule reliability in large academic settings. Additionally, most models lack the capability to generate adaptive schedules that respond in real time to disruptions and evolving requirements, even though mobility conditions and infrastructure availability can change within hours.
@@
Digital twin technology has recently emerged as a transformative tool for addressing these challenges [ye2023developing, ye2024developing]. A digital twin is a dynamic, real-time virtual representation of a physical system that seamlessly integrates spatial-temporal data with operational constraints, offering a powerful platform for simulation, monitoring, and optimization. While digital twins have been widely adopted in manufacturing to enhance production scheduling and operational efficiency [tao2022digital, zhang2021digital, dasgupta2021transportation], their application in urban systems and academic environments is still emerging. For campus operations, a digital twin provides the missing connective tissue between scheduling decisions and human mobility realities: it can encode building layouts, pathway networks, vertical circulation, and evolving conditions (e.g., closures or route changes), enabling timetables to be evaluated and refined based on realistic movement costs and constraints. Recent work in academic scheduling demonstrates the potential of digital twins to adapt dynamically to changes such as fluctuating enrollments, evolving classroom availability, and infrastructure disruptions [ye2023developing]. By merging diverse datasets—including classroom attributes, transportation networks, and user feedback—digital twin technology enables a holistic approach to optimizing campus operations.
Figure 1: Texas A&M Digital Twin.
@@
In this study, we leverage the Texas A&M Campus Digital Twin [ye2024developing, gong2025integrating], as shown in Figure 1, to bridge the identified research gaps through a human-mobility-first perspective. We propose an iterative optimization framework for course timetabling that integrates spatial-temporal correlations, multi-modal transportation considerations, and user-centric mobility metrics. A composite scoring function evaluates course assignments based on normalized measures of classroom occupancy, travel distance, travel time, and vertical transitions between floors, thereby explicitly quantifying mobility burden alongside resource efficiency. The iterative optimization process then generates multiple feasible alternatives for bottleneck courses—those inducing excessive mobility costs or poor utilization—supporting targeted adjustments and dynamic responsiveness. By incorporating real-time feedback and leveraging digital twin capabilities, the proposed framework offers a robust and scalable solution for complex, dynamic scheduling requirements in large academic environments, while directly improving the lived mobility experience of campus users.
@@
This research makes several key contributions. First, it reframes course scheduling as a human mobility-aware resource allocation problem and integrates spatial-temporal factors and multi-modal transportation options into the timetabling process, addressing a critical gap in existing methodologies. Second, it demonstrates the value of an iterative, feedback-driven optimization process for dynamically adapting to evolving constraints and mobility conditions. Third, it highlights the broader applicability of this approach beyond academic scheduling, extending to urban resource management and public facility planning where human movement and accessibility are central concerns. By emphasizing adaptability, transparency, and human-centric mobility outcomes, our work provides a comprehensive framework for addressing the complexities of modern timetabling.
@@ 2.1 Problem Settings
Large-scale course timetabling at universities such as Texas A&M University involves the systematic assignment of a set of courses $C=\{c_{i}\mid i=1,\ldots,N\}$ to a network of classrooms $R=\{r_{j}\mid j=1,\ldots,M\}$ across a series of time slots $T=\{t_{k}\mid k=1,\ldots,K\}$ for a student population $V=\{v_{l}\mid l=1,\ldots,L\}$. The objective is to construct a timetable that not only satisfies stringent institutional policies but also aligns with user preferences and optimizes resource utilization. The complexity arises from the need to meet both hard constraints (e.g., classroom capacity, conflict avoidance, and travel feasibility) and soft constraints (e.g., faculty preferences, degree requirements, and inter-departmental coordination). Moreover, spatial-temporal factors—such as building locations, travel routes, and transition times—introduce additional challenges. This multifaceted problem thus requires a comprehensive framework that assigns courses while simultaneously optimizing the ease of student movement across campus.
@@ 2.2 Variables and Data Representation
[…]
@@
spatial-temporal Data
@@
[…]
@@ 2.3 Objective Function
$\text{Maximize }Z=\alpha_{1}\bar{S}_{O}+\alpha_{2}\bar{S}_{D}+\alpha_{3}\bar{S}_{T}+\alpha_{4}\bar{S}_{F},$ (10)
@@ 2.6 Complexity and Proposed Solution
The integration of numerous hard and soft constraints renders the course timetabling problem a high-dimensional, combinatorial challenge. Traditional static optimization methods often struggle to capture the dynamic, user-centric aspects inherent in such problems. To address these challenges, we propose an iterative recommendation framework that incorporates real-time feedback, detailed spatial-temporal analysis, and dynamic scoring adjustments. This adaptive approach is designed to converge to a high-quality timetable that effectively harmonizes institutional objectives with user satisfaction, thus achieving enhanced adaptability, robustness, and operational efficiency.
@@ 3.3 System Highlights
The proposed recommendation system leverages both spatial-temporal factors and multi-modal transportation data to produce high-quality timetabling solutions. Two key aspects of the system are highlighted below, each incorporating quantitative measures to enhance decision-making.
@@ 4.3 Metric-Based Evaluation
Overall, these metric-based assessments demonstrate the system’s capacity to generate practical, data-driven recommendations that support the twin goals of flexibility and user-centered optimization. By quantifying changes in travel distance, travel time, occupancy, and floor transitions, the framework ensures that proposed reassignments systematically address previously identified spatial-temporal bottlenecks, thereby improving the robustness and efficiency of the course timetable.
@@ 4.4 Comparison of Original and Optimized Schedules
In the histograms on the left, the optimized schedule shows a pronounced shift toward lower travel distances and shorter travel times. The reduction in outliers—students or faculty facing especially long commutes—highlights the system’s effectiveness in targeting spatial-temporal bottlenecks identified during earlier analyses. Additionally, the histogram of floors traveled reveals fewer high-floor transitions in the optimized schedule, suggesting that reassignments have reduced vertical movements between classes. Meanwhile, the occupancy rate distribution indicates a slightly more balanced utilization of classrooms.
@@
Collectively, these results demonstrate how iterative feedback loops—guided by the composite scoring function—can systematically refine scheduling decisions. By quantifying improvements in travel distance, travel time, floors traveled, and occupancy, the framework ensures that adjustments made in each iteration yield tangible benefits. The final outcome is a more user-centric timetable that balances resource utilization with spatial-temporal feasibility, reflecting the adaptive nature of the proposed methodology.
@@ 5 Discussion
A key insight from these experiments was the reduction in travel-related inefficiencies for students and faculty. By incorporating spatial-temporal data, the framework pinpointed and remedied courses that contributed disproportionately to lengthy transitions between classes. In practical terms, this meant fewer instances of extended walks or rushed travel times, which in turn fostered an improved sense of schedule convenience among users. Furthermore, the system’s iterative approach enabled it to react quickly to unexpected events by refining course assignments incrementally, rather than recalculating the entire timetable. Another notable outcome was the framework’s capacity to optimize classroom utilization while respecting critical constraints such as room capacity and scheduling conflicts. By reallocating courses to underused spaces during busy periods, the system minimized congestion without requiring a full overhaul of existing assignments. This flexibility illustrates the framework’s user-centric ethos, wherein data-driven insights guide reassignments that benefit both institutional objectives and individual needs.
@@
Overall, these findings suggest that the proposed recommendation-based framework offers a robust and scalable solution to the complex challenge of course timetabling. By coupling spatial-temporal analytics with iterative refinement and user-centric design, the system not only enhances operational efficiency but also aligns scheduling decisions with the evolving needs of a diverse campus community.
@@ 6 Conclusion
This study introduced an adaptive, recommendation-based framework for course timetabling underpinned by the Texas A&M Campus Digital Twin, with a central emphasis on human mobility. Moving beyond one-shot, static scheduling, the proposed approach reframes the timetable as a dynamic mobility-shaping policy that must remain feasible and useful as constraints, preferences, and operational conditions evolve. In this view, each assignment of a course to a room and time slot is also a decision about movement: it structures when and where students and instructors travel, how they interact with indoor-outdoor infrastructure, and how mobility burdens accumulate across a day. By combining collaborative and content-based filtering with an iterative feedback loop, the framework continuously generates and refines ranked scheduling alternatives, enabling decision-makers to compare mobility-resource trade-offs rather than relying on a single rigid solution. The digital twin serves as the computational substrate that turns mobility from an implicit side effect into an explicit, measurable objective through realistic spatial-temporal costs, supporting responsive scheduling while preserving transparency in how mobility outcomes and institutional constraints are balanced.
Experimental results demonstrate that the framework can systematically reduce mobility friction, lowering travel distance, travel time, and vertical transition burdens, while improving classroom utilization and overall timetable quality. Through bottleneck-oriented refinement, the system identifies courses and transitions that disproportionately amplify movement costs (e.g., long cross-campus transfers, repeated vertical changes, or tightly constrained back-to-back walks) and targets them for localized adjustments, avoiding unnecessary global reshuffling. The resulting schedules exhibit fewer extreme outliers in travel burden, improved feasibility and reliability of consecutive-class transitions, and more balanced occupancy patterns across rooms and time windows. Importantly, these gains are achieved while respecting hard requirements such as capacity and conflict avoidance, showing that mobility-centered objectives and resource efficiency can be optimized within a unified decision process. Beyond quantitative improvements, the framework provides a practical pathway for administrators to operationalize human-centered scheduling by making mobility impacts auditable and comparable across candidate plans.
Despite these strengths, the framework relies on high-quality, comprehensive digital twin data, underscoring the importance of robust integration and validation. Future work can extend the framework with group-wise human movement and interaction-aware relational reasoning to optimize collective congestion dynamics, not only individual travel costs [zhang2024spatial, wu2025hypergraph]. LLM-enabled explainability can translate mobility analytics into stakeholder-facing rationales for proposed changes [liang2024exploring, gong2024mobility, wu2025v2x], while agentic scheduling systems can coordinate specialized agents (mobility, capacity, accessibility, policy) to negotiate and verify timetable updates within the digital twin [du2025cams, zhang2024agentic]. Finally, evolving toward a closed-loop “sense–predict–simulate–act” digital twin pipeline with real-time sensing and predictive models can enable proactive, uncertainty-aware scheduling that incorporates broader objectives such as equity, safety, and sustainability [fan2022integrating, irfan2024toward, wu2025digital].